2/7+4/21+1/3 please help i need to show work also i need to know how to do this problem?
Answer by jim_thompson5910(35256)

In order to add fractions, the denominators must all be the same. So let's get all the denominators equal to the LCD 21.


2/7+4/21+1/3 = (2/7)*(3/3)+4/21+(1/3)*(7/7) = 6/21+4/21+7/21 = (6+4+7)/21 = 17/21


So 2/7+4/21+1/3 = 17/21


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim